Let v=(1,-4,12) and w=(3,5,-1) be vectors.  

What is 2v-3w?
Answer #1
Given
v=(1,-4,12) and w=(3,5,-1)

2v = (2, -8, 24) 
3w = (9, 15, -3)

2v-3w 
= (2-9 -8-15 24+3)
= (-7 -23 27)

(-7 -23 27)